OpenStudy (anonymous):

What is the simplified form of a over 3 all over 6 over 3 a ?

OpenStudy (anonymous):

@zepdrix help please..?

zepdrix (zepdrix):

\[\Large\rm \frac{\left(\frac{a}{3}\right)}{\left(\frac{6}{3a}\right)}\]Did I interpret that correctly? We need to simplify this?

OpenStudy (anonymous):

Yes, that's right.

zepdrix (zepdrix):

When we divide fractions, we can rewrite it as multiplication by flipping the bottom fraction. Example:\[\Large\rm \frac{\left(\frac{e}{f}\right)}{\left(\frac{g}{h}\right)}\quad=\quad \frac{e}{f}\cdot \frac{h}{g}\]And then do normal fraction multiplication, top with top, bottom with bottom,\[\Large\rm \frac{e\cdot h}{f\cdot g}\]And simplify if possible.

OpenStudy (anonymous):

\[a/3 \times 3a/6?\] That would equal 3a^2/18?

zepdrix (zepdrix):

Mmm good! Now it looks like the top and bottom are both divisible by 3. Can we cancel out a 3 somehow?

OpenStudy (anonymous):

Yea. a^2/6

OpenStudy (anonymous):

Okay. Thanks for the help zepdrix!

zepdrix (zepdrix):

yay good job \c:/